Given the following data, perform the appropriate calculations and evaluate the values related to oxygen content and oxygen consumption. PaO2 = 65 mm Hg

Hb = 10 g/dl

SaO2 = 91%

PVO2 = 30 mm Hg

SVO2 = 50%

cardiac output = 16 l/min

FOR EACH QUESTION FOR THIS PROBLEM ANSWER EITHER A. Correct B. Incorrrect.

SHOW FORMULAS OR EXPLANATION FOR EACH QUESTION AS TO WHY IT WAS EITHER CORRECT OR INCORRECT.

CaO2 = 12.39cc O2/dl Correct or Incorrect

a-v DO2 = 5.6 cc O2/dl Correct or Incorrect

This patient is at rest Correct or Incorrect

VO2 = 89.6 cc/min Correct or Incorrect
Answer #1

ANSWER

* CaO2 = 12.39cc O2/dl is (B. Incorrect.)

            Because,

Formula for calculate CaO2 = (Hb * 1.39 * SaO2) + (PaO2 * 0.003)

                                                    Hb = 10 g/dl

                                                  SaO2 = 91% (0.91 gm)

                                                    PaO2 = 65 mm Hg

                                                  CaO2 = (10 * 1.39 *0.91) + (65 * 0.003)

                                                             = 12.65 + 0.195

                                                             = 12.84 cc O2/dl

     * a-v DO2 = 5.6 cc O2/dl is (B. Incorrect.)

              Because,

             Formula = CaO2 - CvO2

                      CaO2    = 12.84cc/dl.

                      SvO2   = 55% (0.55 gm)

                       CvO2      = (Hb * 1.39 * SvO2) +( PaO2 * 0.003)

                                       = ( 10 * 1.39 * 0.55) + (65 * 0.003)

                                       = 7.645 + 0.195

                                       = 7.84cc/dl

                      a - v DO2 = 12.84 - 7.84

                                       = 5 cc/dl.

     * This patient is at rest (A.Correct.)

                    * Measuring arteriovenous diffrence(a-v DO2) help to understand the activity of patient

                      if the patient is in resting position 5ml in every 100ml of blood is extracted and

                    make a arteriovenous oxygen difference(a -v DO2) of 5ml.

      * VO2 = 89.6 cc/ min is (B.incorrect)

              Formula = Cardiac output * (CaO2 - CvO2)

                      Cardiac output = 16L /min.

   = 16 * 5

   = 80 cc/ min